AN ANALYTIC POTENTIAL ENERGY CURVE FOR THE a^{3}Σ^{+} STATE OF KLi, DERIVED FROM OB-SERVATIONS OF THE UPPER VIBRATIONAL LEVELS ONLY

\underline{HOUSSAM SALAMI}, AMANDA J. ROSS, PATRICK CROZET, Université Lyon 1; CNRS; LASIM UMR 5579, 43 Bd du 11 novembre 1918, F-69622 Villeurbanne, France; ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada; PAWEL KOWALCZYK, Institute of Experimental Physics, Warsaw University, ul. Hoża 69, 00681 Warsaw, Poland; WLODZIMIERZ JASTRZEBSKI, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02668 Warsaw, Poland.

Fourier transform resolved fluorescence spectra of the ^{39}K^{6}Li molecule have been recorded in the near-infrared, following laser excitation (Coumarin 6 dye) to high vibrational levels of the B^{1}Π state. The FT spectra show transitions to high vibrational levels of both the X^{1}Σ^{+} and a^{3}Σ^{+} electronic states. These include 147 transitions into six vibrational levels of the a^{3}Σ^{+} state, which lie between 7 and 88 cm^{−1} below the dissociation asymptote. Unfortunately, their energies span less than 30% of the total ground-triplet-state well depth, and even vibrational assignment is not immediately obvious. However, fitting those data to eigenvalues of analytical model potential functions whose outer limbs incorporate the theoretically predicted long-range form, V (R) ≃ D − C_{6}/R^{6} − C_{8}/R^{8} , yields complete, plausible potential curves for this state. The best fits converge to remarkably similar solutions which strongly suggest that the lowest observed level is v=5. Equilibrium parameters from these fits indicate D_{e} = 287(±4) cm^{−1} and R_{e} = 4.99(±0.09) Å for the a^{3}Σ^{+} state of KLi, with ω_{e} = 47.3(±1.5) and 44.2(±1.5) cm^{−1} for ^{39}K^{6}Li and ^{39}K^{7}Li, respectively. Properties of the resulting potential are compared with those of an ab initio potential and with those of the analogous states of Li_{2}, K_{2}, Na_{2} and NaK. The optimized potential curve predicts that the last bound level of both isotopologues lies less than 0.002 cm^{−1} below the atomic asymptote, but since these binding energies are much smaller than the current uncertainty in the ground state dissociation energy (±0.07 cm^{−1}), they remain subject to caution.